hello there,
i'm looking for help solving this problem:
For the function f(x) = 3x^3 + 5x, give a formula for the following functions.
f(-x) = ?
-f(x) = ?
thank you very much, i appreciate your expertise.

Answer by dkppathak(439)

For the function f(x) = 3x^3 + 5x, give a formula for the following functions.
f(-x) = ?
-f(x) = ?
solution
f(x) = 3x^3 + 5x
f(-x) = 3(-x)^3+5(-x)
you have not asked this
f-(-x)=f(x) as both negative convert to positive
f-(-x)= f(x)=3x^3 + 5x,
=3*-x^3-5x=-3^3-5x both term convert to negative as index (power on variable ) is odd
-f(x) = -(3x^3 + 5x)=-3x^3 -5x here also both term convert to negative due to change 0f sign to entire function . it is not always f(-x) = -f(x)
f-(-x)=3x^3 + 5x,